OpenStudy (anonymous):

i'm a little confused here, four different colored flags can be hung in a row to make coded signals. How many flags can be made if a signal consists of the display of one or more flags? I don't know how the answer became 64.

OpenStudy (anonymous):

one flag, 4 ways two flags \(4\times 3=12\) ways three flags \(4\times 3\times 2=24\) ways four flags ...

OpenStudy (anonymous):

i assume it meant "how many SIGNALS...." not how many flags...

OpenStudy (anonymous):

i think so.

OpenStudy (anonymous):

i used permutation and i got 64.

OpenStudy (anonymous):

add em up \[4+12+24+24\]

OpenStudy (anonymous):

ok, thank you.